Rhodes Trucking is considering investing in a new $3.7 million project that will increase net income by 2.7 percent.  This project will be completely funded by issuing new equity shares.  Currently, the firm has 647,400 shares of stock outstanding with a market price of $41 per share.  The current earnings per share of $3.02.  What will the earnings per share be if the project is implemented?
Answer #1

Current Earnings = 3.02 per share * 647400 shares

= 1,955,148  

Revised Earnings. = 1,955,148 * (1+0.027) = $ 2,007,936.996

Existing no of share = 647400 shares

New no of shares issued = 3,700,000 / 41 = 90243.90 OR 90244

Revised no of shares = 647400 + 90244

= 737,644 shares

Revised EPS = Revised Earnings / Revised no of shares

=  2,007,936.996 / 737,644

= $ 2.72 / share

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.